# Machine-learned RG-improved gauge actions and classically perfect gradient flows

### v1

@@ Abstract
In this Letter, we employ an alternative approach, based on fixed-point (FP) lattice actions [15] which are implicitly and nonperturbatively defined using the properties of Wilson’s renormalization group (RG) for asymptotically free theories [16, 17, 18]. It is used to extract continuum quantities from coarse lattices with significantly reduced lattice artifacts. We demonstrate the effectiveness of the FP action in MC simulations of four-dimensional SU(3) gauge-field theory, parameterized using a convolutional neural network with exact gauge invariance [19]. This result is made possible by our new theoretical insight, presented here, that FP actions yield a classically perfect gradient flow, i.e., without tree-level lattice artifacts to all orders in the lattice spacing. This result also builds on the success of our recent work [20, 21, 22], where we demonstrated that ML-based parameterizations can accurately capture the FP action across a wide range of gauge-field fluctuations—from very fine to very coarse lattice spacings. The resulting parameterization achieves an average relative error in the action value of less than $0.2\%$, representing an improvement by over an order of magnitude compared to previous approaches developed earlier [23, 24, 25]. The ML framework also enables accurate parameterization of the derivatives of the FP action w.r.t. the gauge fields, which are crucial for both efficient network training and the practical implementations of MC algorithms. The FP action can be generally used in any lattice-QCD study and provides a new way to reduce discretization artifacts on coarse lattices, making it broadly useful.
@@
[…]
@@
[…]
@@
[…]
@@
where
$C(a^{2}/t)=\frac{64\pi^{2}t^{2}}{3}\times\\ \int_{-\pi/a}^{\pi/a}\frac{d^{4}p}{(2\pi)^{4}}{\rm Tr}\left[e^{-t({\cal S}^{f}+{\cal G})}({\cal S}^{g}+{\cal G})^{-1}e^{-t({\cal S}^{f}+{\cal G})}{\cal S}^{e}\right]$ (7)
contains the tree-level lattice artifacts as deviations from 1. Here, the action density in momentum space to quadratic order in the fields is $A_{\mu}(p){\cal S}_{\mu\nu}(p)A_{\nu}(-p)$, and ${\cal S}^{e},{\cal S}^{f},{\cal S}^{g}$ correspond to the (possibly different) choices of the discretized action for the action-density observable $E$, for the GF, and for the MC simulation, respectively, while ${\cal G}$ is a gauge fixing term. For any choice of action discretizations, the value of $C(a^{2}/t)$ can be calculated and used to define a tree-level improved coupling at finite lattice spacing [34] (see also [35]). Using the same lattice action for all three gives
@@
Machine learning architecture — The practical limitations in using the FP action in MC simulations arise from two factors: the size of quantum lattice artifacts, and the inherent imperfections in the fixed-point parameterization for four-dimensional SU(3) gauge theory. The latter is addressed here using a gauge-equivariant convolutional neural network [19] and ML techniques. The ML architecture designed for the FP action is described in detail in [21]. In brief, the main building block of the neural network is a bilinear convolution which computes various products of Wilson loops, starting from $1\times 1$ plaquettes and the links $U_{n,\mu}$ of a gauge-field configuration as input. Gauge covariance is exactly maintained throughout by accounting for parallel transport in the convolution. As the network depth increases, larger and more complex closed loops are generated. The final layer is a trace giving the action value as output. The network respects translational symmetry and is (ultra-)local, which allows it to be used on any lattice size. Additionally, our architecture is set up to reproduce a perturbatively defined FP action for smooth fields [24] in order to guarantee the correct continuum limit.
@@
The goal of training a model is to approximate the action values according to the FP equation in Eq.(3). A crucial element is that the exact derivatives of ${\cal A}^{\rm FP}$ with respect to gauge links $U_{n,\mu}$ can be easily obtained from the FP equation, with each gauge configuration giving 8×4×8$\times$ 4$\times$8 × 4 ×(volume) data for learning. In the network, the derivatives can be calculated exactly via backpropagation. The loss function for training hence contains both action values and derivatives. The best architecture has been obtained from a hyperparameter scan and consists of three gauge-covariant convolutional layers. It has been trained on equilibrated configurations for a variety of lattice sizes and a large range of coupling parameters from very smooth to coarse gauge-field fluctuations, and has further been improved by finetuning on instanton configurations. Our best model reaches an average relative error in the action value of less than 0.2% with the exact derivatives more accurately reproduced than in previous parameterizations [25] across all lattice spacings. The fidelity of the model for very smooth configurations is particularly important for the GF because the gauge-field fluctuations become very smooth at large flow time.
@@
Fixed-point action simulations — The availability of accurate gauge-field derivatives from our ML parameterization model makes the Hybrid Monte Carlo (HMC) algorithm a suitable choice for simulations, where artificial conjugate momenta are introduced to update the gauge links through molecular dynamics [36, 37]. The simulations can be accelerated with a coarser choice of time step when solving the equations of motion [38] using in particular the 4MN4FP variant of Omelyan integration scheme [39]. The GF for gauge links $dU_{\mu}/dt=-i(\delta{\cal A}^{\rm FP}/\delta U_{\mu})U_{\mu}$ is solved using the third-order Runge-Kutta integrator [30]. The ML network is memory intensive; the largest lattice volumes we simulate are $18^{4}$, for which we can generate a new HMC trajectory in $\sim 4$ minutes on one NVidia A100 SXM6 64GB GPU. With the FP action, a smaller lattice volume is compensated by coarser lattice spacing to yield near-continuum-physics results in large enough physical volume. As a consistency check of continuum predictions with the FP action, we also simulate four-dimensional SU(3) gauge theory using either Wilson or tree-level Symanzik improved gauge actions, measuring with Wilson GF, using the publicly available openqcd package [40].
@@
Scaling tests — The GF provides several physical quantities that can be extracted. For example, a reference scale $t_{c}/a^{2}$ can be set through the condition $t_{c}^{2}\langle E(t_{c})\rangle=c$, where a common choice is $c=0.3$ [30]. An alternative scale $w_{d}/a$ is set through the derivative condition
@@
where $d=0.3$ is also frequently used [41]. At finite lattice spacing, the values of $t_{c}/a^{2}$ and $w_{d}/a$ are determined by the choice of bare coupling $6/g_{0}^{2}$ at which MC simulations are performed, as well as the action discretization used in the GF and the action density. Ratios of scales, e.g., $t_{c}/w_{d}^{2}$ or $t_{d}/t_{c}$, are physical properties, which have a well-defined value in the continuum limit independent of the discretizations employed. The necessary consistency in the continuum limit is a consequence of universality and provides a stringent test of the FP simulation results.
@@
The FP data for $t_{0.3}/w_{0.3}^{2}$ and the $\beta$-function show mild variation with lattice spacing, with artifacts smaller than $1\%$ at $a\simeq 0.14$ fm, using $\sqrt{t_{0.3}}=0.1679\,\mathrm{fm}$ [46] to convert from lattice to physical scales. The ratio $t_{0.5}/t_{0.3}$ has essentially no lattice-spacing dependence below $a\lesssim 0.14$ fm using the FP action. By comparison, discretization errors with the Wilson or Symanzik action and either plaquette or clover version of the action density $E$ are much larger, hence finer lattice spacings and larger lattice volumes are necessary to reach a scaling window of ${\cal O}(a^{2})$ behavior while simultaneously keeping finite-volume effects small.
@@ Acknowledgements.
K.H. wishes to thank the AEC and ITP at the University of Bern for their support. U.W. acknowledges funding from the Swiss National Science Foundation (SNSF) project No. 200020_208222. A.I. and D.M. acknowledge funding from the Austrian Science Fund (FWF) projects P 32446, P 34455, P 34764 and PAT3667424. K.H. acknowledges funding from the US National Science Foundation under grants 2014150 and 2412320. The computational results presented have been achieved in part using the Vienna Scientific Cluster (VSC) and LEONARDO at CINECA, Italy, via an AURELEO (Austrian Users at LEONARDO supercomputer) project, and UBELIX (http://www.id.unibe.ch/hpc), the HPC cluster at the University of Bern.

### v2

@@ Abstract
In this Letter, we employ an alternative approach, based on fixed-point (FP) lattice actions [29] which are implicitly and nonperturbatively defined using the properties of Wilson’s renormalization group (RG) for asymptotically free theories [55, 57, 56]. It is used to extract continuum quantities from coarse lattices with significantly reduced lattice artifacts. We demonstrate the effectiveness of the FP action in MC simulations of four-dimensional SU(3) gauge-field theory, parameterized using a convolutional neural network with exact gauge invariance [18]. This result is made possible by our new theoretical insight, presented here, that FP actions yield a classically perfect gradient flow, i.e., without tree-level lattice artifacts to all orders in the lattice spacing. This result also builds on the success of our recent work [54, 34, 33], where we demonstrated that ML-based parameterizations can accurately capture the FP action across a wide range of gauge-field fluctuations—from very fine to very coarse lattice spacings. The resulting parameterization achieves an average relative error in the action value of less than $0.2\%$, representing an improvement by over an order of magnitude compared to previous approaches developed earlier [15, 9, 43]. The ML framework also enables accurate parameterization of the derivatives of the FP action w.r.t. the gauge fields, which are crucial for both efficient network training and the practical implementations of MC algorithms. It is the first successful application of a machine-learning technique to the simulation of a four-dimensional gauge-field theory which scales to sufficiently large lattice volumes and small lattice spacings to allow for controlled continuum limits of physically relevant quantities. As such it is of general interest both for the broad physics community as well as for the machine-learning community. The machine-learned FP action can be generally used in any lattice-QCD study and provides a new way to reduce discretization artifacts on coarse lattices, making it broadly useful.
Apart from the machine-learning aspect, the classically perfect properties of the FP gradient flow (GF) are of general interest, because it finds applications in many physical fields, e.g., statistical mechanics, fluid dynamics, diffusion processes, optimization problems, and image reconstruction.
@@
In continuum language, for an action $\cal A$} the GF evolves the non-Abelian gauge field $A_{\mu}(t)$ along a fictitious flow time $t$ according to [40, 39, 41, 37]
@@
where the starting value $A_{\mu}(t=0)$ is the original gauge field and the action is ${\cal A}=(1/g^{2})\int d^{4}x(-\frac{1}{2}{\rm Tr}F_{\mu\nu}F_{\mu\nu})$, with field strength $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}+[A_{\mu},A_{\nu}]$ and gauge coupling $g$. The action density $E(t)=-\frac{1}{2}{\rm Tr}F_{\mu\nu}F_{\mu\nu}$ at flow time $t$ is a renormalized quantity, and for SU($N$) gauge theory it can be connected perturbatively to a renormalized coupling $g$ in a given scheme, e.g., the $\overline{\rm MS}$ scheme, as [39]
@@
where $\langle\cdot\rangle$ denotes the gauge-field average.
On the lattice, the perturbative expansion is in the bare coupling $g_{0}$ with [19]
@@
where {multline} C(a^2/t) = 64 π2t23 ×
∫_-π/a^π/a d4p(2 π)4 Tr[e^-t(S^f+G)(S^g+G)^-1e^-t(S^f+G) S^e] contains the tree-level lattice artifacts as deviations from 1. Here, the action density $\cal S$} is in momentum space to quadratic order in the fields, and ${\cal S}^{e},{\cal S}^{f},{\cal S}^{g}$ correspond to the (possibly different) choices of the discretized action for the action-density observable $E$, for the GF, and for the MC simulation, respectively, while $\cal G$} is a gauge fixing term. For any choice of action discretizations, the value of $C(a^{2}/t)$ can be calculated and used to define a tree-level improved coupling at finite lattice spacing [19] (see also [46]). Using the same lattice action for all three gives
@@
Machine learning architecture — The practical limitations in using the FP action in MC simulations arise from two factors: the size of quantum lattice artifacts, and the inherent imperfections in the fixed-point parameterization for four-dimensional SU(3) gauge theory. The latter is addressed here using a gauge-equivariant convolutional neural network [18] and ML techniques. The ML architecture designed for the FP action is described in detail in [34]. In brief, the main building block of the neural network is a bilinear convolution which computes various products of Wilson loops, starting from $1\times 1$ plaquettes and the links $U_{n,\mu}$ of a gauge-field configuration as input. Gauge covariance is exactly maintained throughout by accounting for parallel transport in the convolution. As the network depth increases, larger and more complex closed loops are generated. The final layer is a trace giving the action value as output. The network respects translational symmetry and is (ultra-)local, which allows it to be used on any lattice size. Additionally, our architecture is set up to reproduce the Yang-Mills action in the naive continuum limit and approximates a perturbatively defined FP action on smooth fields [9], see [34] for further details.
@@
The goal of training a model is to approximate the action values according to the FP equation in Eq.(3). A crucial element is that the exact derivatives of ${\cal A}^{\rm FP}$ with respect to gauge links $U_{n,\mu}$ can be easily obtained from the FP equation, with each gauge configuration giving $8\times 4\times$(volume) data for learning. In the network, the derivatives can be calculated exactly via backpropagation. The loss function for training hence contains both action values and derivatives. The best architecture has been obtained from a hyperparameter scan and consists of three gauge-covariant convolutional layers. It has been trained on equilibrated configurations for a variety of lattice sizes and a large range of coupling parameters from very smooth to coarse gauge-field fluctuations, and has further been improved by finetuning on instanton configurations. Our best model reaches an average relative error in the action value of less than 0.2% with the exact derivatives more accurately reproduced than in previous parameterizations [43] across all lattice spacings. The fidelity of the model for very smooth configurations is particularly important for the GF because the gauge-field fluctuations become very smooth at large flow time.
@@
Fixed-point action simulations — The availability of accurate gauge-field derivatives from our ML parameterization model makes the Hybrid Monte Carlo (HMC) algorithm a suitable choice for simulations, where artificial conjugate momenta are introduced to update the gauge links through molecular dynamics [17, 24]. The simulations can be accelerated with a coarser choice of time step when solving the equations of motion [45] using in particular the 4MN4FP variant of Omelyan integration scheme [53]. The GF for gauge links $dU_{\mu}/dt=-i(\delta{\cal A}^{\rm FP}/\delta U_{\mu})U_{\mu}$ is solved using the third-order Runge-Kutta integrator [40]. The ML network is memory intensive, but the memory usage only scales linearly with the lattice volume; the largest lattice volumes we simulate are $18^{4}$, for which we can generate a new HMC trajectory in $\sim 4$ minutes on one NVidia A100 SXM6 64GB GPU. With the FP action, a smaller lattice volume is compensated by coarser lattice spacing to yield near-continuum-physics results in large enough physical volume. As a consistency check of continuum predictions with the FP action, we also simulate four-dimensional SU(3) gauge theory with the publicly available openqcd package [60] using either Wilson or tree-level Symanzik improved gauge actions and measuring with Wilson GF. [^1] This generic choice, instead of the alternative schemes in Refs. [19, 46], is justified by the fact that we are first of all interested in the continuum values which are universal and independent of the choice. Moreover, when comparing lattice artifacts, the improvement schemes are highly problem-specific, while the FP approach is general and applies to other observables.
@@
Scaling tests — The GF provides several physically relevant quantities that can be extracted. For example, a reference scale $t_{c}/a^{2}$ can be set through the condition $t_{c}^{2}\langle E(t_{c})\rangle=c$, where a common choice is $c=0.3$ [40]. An alternative scale $w_{d}/a$ is set through the derivative condition
@@
where $d=0.3$ is also frequently used [10]. These scales essentially probe the gauge-field fluctuations at specific distances parameterized by the parameters $c$ and $d$. At finite lattice spacing, the values of $t_{c}/a^{2}$ and $w_{d}/a$ are determined by the choice of bare coupling $6/g_{0}^{2}$ at which MC simulations are performed, as well as the action discretization used in the GF and the action density. Ratios of scales, e.g., $t_{c}/w_{d}^{2}$ or $t_{d}/t_{c}$, are physical properties, which have a well-defined value in the continuum limit independent of the discretizations employed. The necessary consistency in the continuum limit is a consequence of universality and provides a stringent test of the FP simulation results.
@@
The FP data for $t_{0.3}/w_{0.3}^{2}$ and the $\beta$-function show mild variation with lattice spacing, with artifacts smaller than $1\%$ at $a\simeq 0.14$ fm, using $\sqrt{t_{0.3}}=0.1679\,\mathrm{fm}$ [23] to convert from lattice to physical scales. The ratio $t_{0.5}/t_{0.3}$ has essentially no lattice-spacing dependence below $a\lesssim 0.14$ fm using the FP action. By comparison, when discretization errors are larger, finer lattice spacings and larger lattice volumes are necessary to reach a scaling window of ${\cal O}(a^{2})$ behavior while simultaneously keeping finite-volume effects small. [^3] We note that a more meaningful discussion of the relative cost of simulations using standard and FP actions is a delicate and difficult task as it depends heavily on the chosen benchmarks and targets for comparison, e.g., observables, level of lattice artifacts, architecture-specific code optimization, etc., and is hence not attempted here.
@@ Acknowledgements.
K.H. wishes to thank the AEC and ITP at the University of Bern for their support. U.W. acknowledges funding from the Swiss National Science Foundation (SNSF) project No. 200020_208222. A.I. and D.M. acknowledge funding from the Austrian Science Fund (FWF) projects P 32446, P 34455, P 34764 and PAT3667424. K.H. acknowledges funding from the US National Science Foundation under grants 2014150 and 2412320. The computational results presented have been achieved in part using the Vienna Scientific Cluster (VSC) and LEONARDO at CINECA, Italy, via an AURELEO (Austrian Users at LEONARDO supercomputer) project, and UBELIX, the HPC cluster at the University of Bern.